OpenStudy (anonymous):

if f(x)= x+3 / 4 (x+3 over 4), what is the equation for f^–1(x)? a. f^–1(x) = 4(x - 3) b. f^–1(x) = 4x + 3 c. f^–1(x) = 4(x + 3) d. f^–1(x) = 4x - 3

OpenStudy (jhannybean):

\[f(x)=\frac{x+3}{4\left(\dfrac{x+3}{4}\right)}\]Is this your problem?

OpenStudy (jdoe0001):

\(\bf f(x)={\color{blue}{ y}}=\cfrac{{\color{brown}{ x}}+3}{4} \quad \textit{do the switcharoo baby}\quad {\color{brown}{ x}}=\cfrac{{\color{blue}{ y}}+3}{4}\impliedby f^{-1}(x)\)

OpenStudy (anonymous):

\[f(x) = \frac{ x+3 }{ 4 }\]

OpenStudy (campbell_st):

so the function is \[y = \frac{x + 3}{4}\] to find the inverse swap x and y \[x = \frac{y + 3}{4}\] now make y the subject of the equation

OpenStudy (jhannybean):

Oh I simply read it wrong, whoops.